OpenStudy (anonymous):

Lester is constructing two similar boxes shaped as similar rectangular prisms. The larger box has a height of 18 inches and a square base of side length 15 inches. If the smaller box has a height of 12 inches, what is its volume?

OpenStudy (campbell_st):

Large box volume = V = 15^2 x 18 V = 4050 inches^3 Small box since similar ratios are the same for all sides... 12/18 = 2/3 the small box has measurements 2/3 the size of the large box so the base lengths are 2/3 x 15 = 10 small box volume V = 10^2 x 12 V = 1200 inches ^3 the alternative method is (2/3)^3 x volume of the large box... 8/27 x 4050 = 1200 inches^3 since you are finding volume the scaling factor is cubed... if it was area the scaling factor is squared. hope this helps